Program Information
Recent Log Entries
- Vacuum Group Summary 2026-02-18
- The following cavity histories have been updated: 1L20-4
- COMPLETED Measure VIP1L25-1 HVMON logic bit response time with oscilloscope
- COMPLETED Troubleshoot Hall D Tagger dump vacuum power supply
- COMPLETED Manually Tune Separator Cavity 3-2
- QuickPic - Pfwd indication bad on 1L11-3
- CREW CHIEF SIGN IN
- Follow-up Re: Checked UVA RTP Signals
- RF On/Off turned RF On
- Survey and Alignment work summary
- COMPLETED Troubleshoot North Access Cooling Tower Pump Controls
- COMMENT ON C75-05 (2L03) SRF Cryomodule Commissioning
- DecaRads are on 1L05, 1L10, 1L25
- COMMENT ON SL25 and SL26 vacuum controls and cable installation for reworked 2L25/2L26 girders
- COMMENT ON Post SAM FE onset measurements C75 and C100 cryomodules











